OpenStudy (anonymous):

In a certain class, 1/2 of the students are ballers and 1/3 of the ballers are seniors. If 1/2 of the senior ballers have failed the last test, which of the following could be the total number of students in the class? A. 3 B. 9 C. 12 D. 18 E. 21 the questions seems weird.. not sure if it's me, though

OpenStudy (fibonaccichick666):

so first, let's look at how many ballers are in the class. Let us say that there are x ballers in the class. Then \(\frac{1}{2} x\) is the number of students that are ballers in general. Do you follow?

OpenStudy (anonymous):

mhm. then 1/6 will be seniors, correct? is that what you're getting at?

OpenStudy (fibonaccichick666):

yep then what is the maximum fraction of the class that was senior that could have failed?

OpenStudy (anonymous):

then 1/12 will be the ones that failed and since you can't have 1/12th of any other number other than 12, answer is C. YOU, MY FRIEND, ARE A GENIUS! :D

OpenStudy (fibonaccichick666):

wait wait no

OpenStudy (anonymous):

*suspense*

OpenStudy (fibonaccichick666):

you have to do the other operators and work backwards still

OpenStudy (anonymous):

papoi? o.0

OpenStudy (fibonaccichick666):

Ok I don't think any of them are right because you cannot take half and then get an appropriate answer for further division

OpenStudy (anonymous):

but 1/12th of 12 is one. so 1 senior baller failed. 6 of them are ballers. 2 of them are seniors D:

OpenStudy (fibonaccichick666):

ignore that last 12 half is 6 1/3 of 6 is 2 1/2 is 1

OpenStudy (fibonaccichick666):

just read wrong I oops

OpenStudy (fibonaccichick666):

yea you're right lol

OpenStudy (anonymous):

thank you very much! :) I'm trying to help my brother with SAT stuff and this question was getting on my nerves XD

OpenStudy (fibonaccichick666):

lol yea, it's fun. The original method worked, but I ran it forwards instead of backwards so I took \(\frac{1}{12}\) instead of \(\frac{1}{2}\) haha